OpenStudy (shiburinga):

Mathematically speaking what is 9*50/100-90+1000=

OpenStudy (shiburinga):

YEY Mhm no more bucks :)

pooja195 (pooja195):

Use PEMDAS

OpenStudy (alexandervonhumboldt2):

\[\frac{ 9*50 }{ 100-90+1000 }\]

OpenStudy (alexandervonhumboldt2):

first muuplty

OpenStudy (alexandervonhumboldt2):

9*50=?

pooja195 (pooja195):

@AlexandervonHumboldt2 It would be best if you let them set it up dont rush it.

OpenStudy (alexandervonhumboldt2):

@pooja195 please stop, this is my first qh question i asnwer .

OpenStudy (alexandervonhumboldt2):

@ShiBuringa what is 9*50=??

OpenStudy (bibby):

@pooja195 please don't spam questions

OpenStudy (shiburinga):

Lol guys it was a question so I could give the bucks :]

OpenStudy (alexandervonhumboldt2):

gimme me bucks hahhaa

OpenStudy (shiburinga):

I think I already did

OpenStudy (alexandervonhumboldt2):

you have t rate me

pooja195 (pooja195):

You don't ask to be rated. The asker will rate you when they are satisfied with your help.

OpenStudy (alexandervonhumboldt2):

i just say how to reward, i dont ask to be rated

TheSmartOne (thesmartone):

Why are you guys asking to be rated before even answering the question? ._.

OpenStudy (alexandervonhumboldt2):

i'm not asking

TheSmartOne (thesmartone):

First, we need to get the exact question. @ShiBuringa could you write the question using the equation tool or the draw tool?

OpenStudy (alexandervonhumboldt2):

ShiBuringa Best Response Undo Medals 3 Lol guys it was a question so I could give the bucks :] are you even reading what she said tso?

TheSmartOne (thesmartone):

Because your question could either be: a) \(\Large\frac{ 9\cdot50 }{ 100-90+1000 }\) b) \(\Large\frac{ 9\cdot50 }{ 100-90 }+1000\) c) \(\Large\frac{ 9\cdot50 }{ 100}-90 + 1000\)

OpenStudy (shiburinga):

Ok am back was doing dishes

TheSmartOne (thesmartone):

@AlexandervonHumboldt2 Nonetheless, we need to get the correct question to be properly able to assist.

OpenStudy (shiburinga):

@TheSmartOne I made this question to rate my fellow Qualified Helpers, because they have helped me well. :)

TheSmartOne (thesmartone):

[No comment]

OpenStudy (shiburinga):

And at the time I didn;'t know how to rate them xP

OpenStudy (shiburinga):

How do I have three medals .-.

TheSmartOne (thesmartone):

4 medals* ;)

OpenStudy (shiburinga):

Now 4 .-.

OpenStudy (shiburinga):

lol:P

OpenStudy (shiburinga):

AW GUYS WHY OS WHY

OpenStudy (shiburinga):

NOO

OpenStudy (shiburinga):

*cri*

OpenStudy (shiburinga):

I can only rate one >.<

OpenStudy (shiburinga):

I can only rate one >.<

TheSmartOne (thesmartone):

Rate whoever you want :)

jagr2713 (jagr2713):

:D Glad to see the Ob's going to great use :D

hartnn (hartnn):

serious answer to a non-serious question :P first priority : PEDMAS and then left to right. 9*50/100-90+1000 = 450/100-90+1000 = 0.45 - 90 + 1000 (now the order doesn't matter as all operations are of same priority) = 910.45

TheSmartOne (thesmartone):

@hartnn but we need the correct question. You gave the answer to a question that could have 3 different possibilities, which I listed above :P

OpenStudy (shiburinga):

Mhm xP

OpenStudy (shiburinga):

Lol! Hartnn... it was serious to a rate a QH. xP

hartnn (hartnn):

TSO, assuming the asker didn't miss any brackets, there could be only one possibility of the question :) Shi, we appreciate your gesture to rate a helper who had helped you earlier. Thanks a lot for that!

OpenStudy (shiburinga):

Np. :]

TheSmartOne (thesmartone):

:P